OpenStudy (anonymous):

A wooden duck with a regular selling price of $109.45 is marked down to $69.50. What is the percent of markdown based on the regular price? (Round to nearest tenth percent.)

hero (hero):

\[\frac{\%\Delta}{100} = \frac{a - b}{a}\] \(\%\Delta\) = percent change a = original amount b = discounted amount

OpenStudy (anonymous):

what is the answer then?

hero (hero):

I'll help you set it up. You can compute it on your own.

hero (hero):

\[\frac{\%\Delta}{100} = \frac{109.45 - 69.50}{109.45}\]

OpenStudy (anonymous):

welcome to openstudy please understand the goal of this website is to help you learn how to do the problems please do not come here expecting us to give you the answer also please read the code of conduct or terms and condition if you get a chance